# pnasresearcharticle

### v6

@@ Abstract
De novo peptide sequencing $|$ Deep learning $|$ Data independent acquisition
[…]
This manuscript was compiled on September 19, 2025 www.pnas.org/cgi/doi/10.1073/pnas.XXXXXXXXXX
[…]
@@
De novo peptide sequencing plays a crucial role in proteomics, enabling the identification of novel peptides, post-translational modifications, and mutations absent from existing protein databases(1)(2)(3)(4). This capability is crucial for personalized immunotherapy, guiding targeted treatments by identifying unique neoantigens. It is also essential for studying species with unsequenced genomes, where database searches are not feasible (5) (6) (7).
@@
Compared to database-driven methods, de novo sequencing offers the advantage of discovering new peptide sequences independently of pre-existing data, enables capturing the full diversity of proteomes, especially in complex and dynamic biological systems (7) (8) (9). In traditional Data-Dependent Acquisition (DDA) methods, where abundant peptides are selected to fragment(10) (11), deep learning-based approaches such as DeepNovo(2), Casanovo(12), PepNet(13), and GraphNovo(14) have significantly improved performance, making them valuable and efficient tools for biological research.
@@
Despite their promising results, traditional DDA methods suffer from limitations such as biased sampling and inconsistent detection of low-abundance peptides, leading to incomplete proteome coverage (11) (15) (16). Data-Independent Acquisition (DIA) (16)(17) addresses these limitations by fragmenting all ionized peptides within a predefined mass range, providing a more comprehensive and unbiased snapshot of the proteome (15).
@@
Data-independent acquisition (DIA) presents unique challenges for peptide sequencing compared to data-dependent acquisition (DDA). Unlike DDA, which typically produces spectra associated with a single precursor ion, DIA generates highly multiplexed spectra in which fragment ions from multiple coeluting peptides coexist (18) (19). This concurrent fragmentation increases spectral complexity and introduces substantial noise, making it difficult to distinguish true signals from background interference. The overlap of fragmentation patterns further complicates fragment-ion assignment, thereby posing significant obstacles for de novo peptide sequencing and requiring more computationally intensive analysis pipelines. Despite these challenges, DIA also provides rich chromatographic information by capturing the temporal profiles of peptide ions. These chromatogram traces, when compared across coeluting peptides, can offer valuable clues for linking fragment ions to their peptide sources, thus partially mitigating the interpretative difficulties inherent in DIA data and offering advantages for downstream peptide identification and quantification.
@@
Recent advances such as DeepNovo-DIA (20), Transformer-DIA (21), and Cascadia (22) have leveraged deep learning to address the unique challenges of de novo sequencing in DIA. These models employ neural architectures to jointly analyze precursor and fragment ions across multiple dimensions, including mass-to-charge ratio ($m/z$), retention time, and intensity, thereby enabling more effective interpretation of highly multiplexed spectra. For example, DeepNovo-DIA integrates Ion-CNN and Spectrum-CNN (23) with a long short-term memory (LSTM) network (24) to capture both the three-dimensional structure of fragment ions and their correlations. Transformer-DIA, in contrast, adopts a Transformer-based encoder–decoder framework (25), autoregressively predicting peptide sequences by generating each amino acid conditioned on the previously decoded subsequence. Cascadia (22) further extends this paradigm by adapting a Transformer model originally designed for DDA data to DIA through the use of large-scale training data and spectrum augmentation strategies. Despite these successes, current approaches remain limited by their tendency to encode only subsets of the available spectral evidence, often neglecting informative signals arising from coeluting peptides, which may reduce sensitivity for detecting challenging peptide sequences.
@@
Figure 1: (a) The model structure of Our entire workflow. On the top is the optimal path task, generating a series of node indices, transformed into the optimal path. The mass values in the optimal path are then translated to the corresponding amino acids when a single match is found. On the button is the sequence generation task. It takes the generated optimal path and outputs the amino acid sequence to replace mass tags. (b) The BERT(26)-like pretrain model (c) An example of a spectrum graph, where the bottom value on each edge represents the mass difference between nodes, encoded by RoPE, and the top value indicates the corresponding amino acid sequence. Only a subset of nodes and edges is plotted for clarity, whereas, in a complete spectrum graph, all possible forward connections would be present.
@@
We propose DIANovo, a framework specifically designed to address the complexities of DIA data introduced by coelution, with its architecture shown in Figure 1(a). Concretely, we first set to tackle the sheer size of highly-multiplexed DIA spectrum. Our encoder encodes the spectrum graph, like GraphNovo(14), as shown in Figure 1(c). To reduce memory consumption, we implements an automated edge generation process, which primarily focuses on learning the relative mass differences between graph nodes while the inter-node amino acid information is learned in a task-specific manner. Moreover, we introduce a coelution-aware pretraining step to incorporate information about coeluting peptides alongside the target peptide, shown in Figure 1(b). This involves pretraining a model to predict ion types from coeluting peptides, with the resulting embeddings used as features in subsequent training stages. The coelution information helps the model better differentiate between signal and noise, leading to more accurate target peptide predictions.
@@
Additionally, our study seeks to assess the realistic performance of DIA de novo sequencing, by reporting amino acid and peptide recalls and precisions compared with DIA-NN (27) search outcomes, considering only peptides unique to the test set. Our findings indicate that de novo peptide sequencing using DIA on older-generation instruments like the Q Exactive(28) or Fusion (29) is suboptimal, yielding relatively lower peptide recall. However, the Orbitrap Astral (30) significantly outpaces older-generation in acquisition speed, enabling the use of narrow-window data-independent acquisition (nDIA). The Astral narrow-window DIA method (30), which offers more consistent fragmentation patterns and fewer missing fragmentations, even in case of high coelution level, yields better results. Our extensive experiments across various datasets, including older-generation and Astral data, highlight the robustness and effectiveness of our proposed method. We demonstrated that our model consistently outperformed the baselines. These findings underscore the potential of our method for robust and accurate de novo peptide sequencing in the challenging DIA setting.
@@
To better understand these challenges, we developed a theoretical framework that explains how the balance between signal enhancement and noise accumulation in different acquisition methods affects de novo peptide sequencing performance. Specifically, we analyze the signal and noise characteristics of DDA, older-generation DIA, and Astral DIA, and their impact on the efficacy of peptide matching algorithms like XCorr.(31) (32) Our model reveals that older-generation DIA introduces substantial noise that outweighs the marginal gain in signal peaks—diminishing de novo sequencing performance—while Astral DIA provides a disproportionate increase in signal peaks despite higher noise levels. This increase effectively enhances peptide identification by improving the confidence of peptide-spectrum matches. This theoretical insight underscores the critical importance of optimizing signal-to-noise profiles in mass spectrometry data to improve de novo sequencing outcomes.
@@ Results
Furthermore, in the comparison of DDA and DIA on the Hela (33) (older-generation) and PXD046453 (30) Hek293T (Astral) dataset, DIA demonstrated advantage in peptide detection when isolation window is small, with our de novo sequencing algorithm identifying additional peptides that DDA missed. These findings provide valuable insights into the proper scenarios for applying DIA de novo sequencing and highlight DIA’s capability to extend identification beyond DDA.
@@
We compared our model’s performance with the baselines across several datasets (34) (35) (33) in Figure 2, where peptide recall is defined as the proportion of peptides whose entire sequences exactly match the database search results, without any mismatches or errors. On most datasets, our model demonstrated significantly better performance than the baselines, however the results are relatively lower across the board compared to DDA levels (2) (14), indicating that there is a significant gap between database and de novo identification performance in older-generation mass spectrometers. Our amino acid recall is on average 60% higher than DeepNovo-DIA, while peptide recall being 53% higher. Comparing to Transformer DIA, our amino acid recall and peptide recall is 32% and 24% higher respectively.
@@
For older-generation instruments, we adopted the Hela dataset (33), analyzing peptides identified from the same biological sample in the DDA and DIA modes with varying DIA isolation windows (5 Th, 10 Th, and 20 Th). In the DDA experiments, we performed a database search using the PEAKS(36) DB search engine and employed PointNovo (37) for de novo peptide sequencing. For DIA, we utilized the DIA-NN(27) search engine for database search and applied our de novo sequencing methods.
@@
In contrast, the performance of the Orbitrap Astral mass spectrometer, using the PXD046453 (30) HEK293T dataset, presents a more consistent scenario (shown in Figure 5) due to the enabled narrow window mode. This dataset provides both DDA data and DIA data acquired at different cycle times (2.5 ms and 3.5 ms), both using a narrow 2 Th isolation window. The Astral data demonstrates that DIA consistently outperforms DDA with the narrow window mode enabled. The DDA de novo sequencing identifies approximately 17,000 peptides, whereas the DIA de novo sequencing detects nearly 32,000 peptides, showing a considerable increase over DDA. The additional 15,000 peptides identified by DIA de novo are likely missed by DDA due to biased sampling. This highlights the key advantage of DIA over DDA—its comprehensive and unbiased sampling of peptides, which is further enhanced by the speed and sensitivity of the Orbitrap Astral. In this case, the smaller isolation window and improved performance of Astral’s DIA mode enable the identification of a broader range of peptides, overcoming the limitations observed with DDA.
@@
Comparison with Cascadia (22)
@@
In database searches, the XCorr score (31) quantifies the similarity between the experimental spectrum and theoretical spectra generated from candidate peptides in the database. A higher XCorr score reflects a better match, leading to increased confidence in peptide identification. This is because the probability of a high-scoring match occurring by chance decreases exponentially with increasing score, enhancing the specificity of the identification. For de novo sequencing, the advantage of a higher XCorr score stems from the intrinsic similarities between database search algorithms and de novo methods. De novo algorithms are heuristic approaches designed to reconstruct peptide sequences directly from spectra without relying on a predefined database. Although de novo methods may employ advanced scoring functions tailored for sequence reconstruction, the XCorr score serves as a valuable lower-bound approximation of their performance. A higher XCorr score implies that the spectrum contains clear and informative fragment ion peaks that can be effectively utilized by de novo algorithms to deduce the peptide sequence.
@@
To validate our theory that signal-noise profile plays a crucial rule in de novo performance, we conducted simulations by performing a grid search through signal peak and noise peak numbers by generating synthetic spectra. Utilizing the approach proposed by Noble et al. (2012)(32), we computed the p-value for a single peptide-spectrum match based on dynamic programming and XCorr scoring. This method allows for the estimation of the statistical significance of the observed XCorr scores, providing insight into the confidence of peptide identifications under varying signal-to-noise conditions.
@@ Discussion
To better understand the statistical behavior of de novo peptide sequencing, we performed controlled simulations to generate synthetic spectra that mimic the signal and noise characteristics of different acquisition methods and test datasets. Signal peaks were sampled from theoretical spectra to represent peptide evidence, while noise peaks were added at levels representative of typical coelution and background noise for each acquisition method. For each simulated spectrum, we computed XCorr scores by cross-correlating the simulated spectra with the corresponding theoretical spectra, followed by significance estimation using a dynamic programming approach for p-value calculation (32). Our analysis further examined the impact of multiple hypothesis testing in de novo sequencing, where the dramatically larger search space compared to database searches necessitates stricter statistical corrections, such as the Šidák adjustment (38). This adjustment increases the stringency of significance thresholds, which in turn reduces peptide recall in de novo sequencing relative to database-driven methods. This theoretical derivation further illustrates that database searches are less impacted by the elevated p-values arising from the degraded signal-to-noise profiles observed when transitioning from DDA to older-generation DIA. This resilience helps database search methods maintain reasonable performance under noisier conditions, whereas de novo sequencing, with its vastly larger hypothesis space, experiences a more pronounced drop in performance.
@@
Code and Data Availability
[…]
@@
[…]
[34]
@@ Appendix A The DIANovo Model
The feature construction for the encoder is designed to capture the chromatogram characteristics in DIA, in order to better distinguish signal peaks from coeluting fragment ions and noise. For each PSM (peptide-spectrum match), we collect five neighboring MS2 spectra along the retention time (RT) dimension, centered around the RT peak. The neighboring spectra provide a chromatogram for each observed peak in the spectrum, and the relationship among these chromatograms improves the robustness of peptide sequence predictions. (27). During preprocessing, peaks across neighboring spectra with similar $m/z$ values are grouped together, enabling the construction of chromatograms for each m/z value. This is achieved by binning each spectrum into fixed-size 0.01 Th intervals, and once the chromatograms are generated through binning, the spectra are reconstructed as a list of (m/z, chromatogram) pairs for neural network processing. The binning approach offers greater computational efficiency than methods such as KD-trees, making it particularly well-suited for the characteristics of our DIA data.
@@
In addition to the primary spectral data, we also included eight supplementary features to enrich the feature set used for encoding, adapted from GraphNovo(14) and Novor(39). These features are derived from each spectrum within the $\pm 50$ Th neighboring window and are designed to capture various statistical and intensity-based characteristics of the spectral peaks.
@@
where offset $\sigma$ depends on the ion type, computed as Supporting Table 1; $c$ is the charge of ion; $m_{proton}$ is the mass of proton and $m_{precursor}$ is the precursor mass (14).
@@
In addition, we adopt the FlashAttention 2 (40) implementation of the attention function, allowing us to process very large DIA spectra, with high computation and memory efficiency.
@@
GraphNovo employs the Graphformer(41) graph neural network to encode the spectrum graph, capturing comprehensive information about the spectrum. However, applying this approach to DIA data is impractical due to the large size of DIA spectra, which results from high levels of coelution. To address this challenge, we replace the memory-intensive node and path encoders used in Graphformer with Rotary Position Embedding (RoPE). This step is also neccessary for the adoption of Flash Attention 2, since it does not allow us to operate directly on the attention matrix, which is required by Graphformer.
RoPE is a position encoding technique that introduces rotational invariance by representing positional information in a circular format(42). In our approach, RoPE encodes the mass differences between graph nodes, allowing the model to learn meaningful mass relationships automatically. This effectively transforms the spectrum graph into a fully-connected graph where edges represent mass differences, enabling the model to identify the most relevant mass differences for the task without the computational overhead of traditional node and path encoders. The adoption of RoPE is advantageous, as it captures the critical information conveyed by the mass difference between pairs of graph nodes, effectively representing the cumulative mass of the amino acids that lie between them.
@@
We adopt a two-stage decoding process, similar to GraphNovo(14) to alleviate the sequence memorization issue during training, where the decoder simply remembers seen training sequences, and fail to generalize on unseen sequences. This approach includes:

### v7

# Disentangling the Complex Multiplexed DIA Spectra in De Novo Peptide Sequencing

@@ Abstract
Keywords: De novo peptide sequencing; Deep learning; Data independent acquisition
1 Introduction
@@
De novo peptide sequencing plays a crucial role in proteomics, enabling the identification of novel peptides, post-translational modifications, and mutations absent from existing protein databases[1][2][3][4]. This capability is crucial for personalized immunotherapy, guiding targeted treatments by identifying unique neoantigens. It is also essential for studying species with unsequenced genomes, where database searches are not feasible [5] [6] [7].
@@
Compared to database-driven methods, de novo sequencing offers the advantage of discovering new peptide sequences independently of pre-existing data, enables capturing the full diversity of proteomes, especially in complex and dynamic biological systems [7] [8] [9]. In traditional Data-Dependent Acquisition (DDA) methods, where abundant peptides are selected to fragment[10] [11], deep learning-based approaches such as DeepNovo[2], Casanovo[12], PepNet[13], and GraphNovo[14] have significantly improved performance, making them valuable and efficient tools for biological research.
@@
Despite their promising results, traditional DDA methods suffer from limitations such as biased sampling and inconsistent detection of low-abundance peptides, leading to incomplete proteome coverage [11] [15] [16]. Data-Independent Acquisition (DIA) [16][17] addresses these limitations by fragmenting all ionized peptides within a predefined mass range, providing a more comprehensive and unbiased snapshot of the proteome [15].
@@
Data-independent acquisition (DIA) presents unique challenges for peptide sequencing compared to data-dependent acquisition (DDA). Unlike DDA, which typically produces spectra associated with a single precursor ion, DIA generates highly multiplexed spectra in which fragment ions from multiple coeluting peptides coexist [18] [19]. This concurrent fragmentation increases spectral complexity and introduces substantial noise, making it difficult to distinguish true signals from background interference. The overlap of fragmentation patterns further complicates fragment-ion assignment, thereby posing significant obstacles for de novo peptide sequencing and requiring more computationally intensive analysis pipelines. Despite these challenges, DIA also provides rich chromatographic information by capturing the temporal profiles of peptide ions. These chromatogram traces, when compared across coeluting peptides, can offer valuable clues for linking fragment ions to their peptide sources, thus partially mitigating the interpretative difficulties inherent in DIA data and offering advantages for downstream peptide identification and quantification.
@@
Recent advances such as DeepNovo-DIA [20], Transformer-DIA [21], and Cascadia [22] have leveraged deep learning to address the unique challenges of de novo sequencing in DIA. These models employ neural architectures to jointly analyze precursor and fragment ions across multiple dimensions, including mass-to-charge ratio ($m/z$), retention time, and intensity, thereby enabling more effective interpretation of highly multiplexed spectra. For example, DeepNovo-DIA integrates Ion-CNN and Spectrum-CNN [23] with a long short-term memory (LSTM) network [24] to capture both the three-dimensional structure of fragment ions and their correlations. Transformer-DIA, in contrast, adopts a Transformer-based encoder–decoder framework [25], autoregressively predicting peptide sequences by generating each amino acid conditioned on the previously decoded subsequence. Cascadia [22] further extends this paradigm by adapting a Transformer model originally designed for DDA data to DIA through the use of large-scale training data and spectrum augmentation strategies. Despite these successes, current approaches remain limited by their tendency to encode only subsets of the available spectral evidence, often neglecting informative signals arising from coeluting peptides, which may reduce sensitivity for detecting challenging peptide sequences.
@@
Figure 1: (a) The model structure of Our entire workflow. On the top is the optimal path task, generating a series of node indices, transformed into the optimal path. The mass values in the optimal path are then translated to the corresponding amino acids when a single match is found. On the button is the sequence generation task. It takes the generated optimal path and outputs the amino acid sequence to replace mass tags. (b) The BERT[26]-like pretrain model (c) An example of a spectrum graph, where the bottom value on each edge represents the mass difference between nodes, encoded by RoPE, and the top value indicates the corresponding amino acid sequence. Only a subset of nodes and edges is plotted for clarity, whereas, in a complete spectrum graph, all possible forward connections would be present.
@@
We propose DIANovo, a framework specifically designed to address the complexities of DIA data introduced by coelution, with its architecture shown in Figure 1(a). Concretely, we first set to tackle the sheer size of highly-multiplexed DIA spectrum. Our encoder encodes the spectrum graph, like GraphNovo[14], as shown in Figure 1(c). To reduce memory consumption, we implements an automated edge generation process, which primarily focuses on learning the relative mass differences between graph nodes while the inter-node amino acid information is learned in a task-specific manner. Moreover, we introduce a coelution-aware pretraining step to incorporate information about coeluting peptides alongside the target peptide, shown in Figure 1(b). This involves pretraining a model to predict ion types from coeluting peptides, with the resulting embeddings used as features in subsequent training stages. The coelution information helps the model better differentiate between signal and noise, leading to more accurate target peptide predictions.
@@
Additionally, our study seeks to assess the realistic performance of DIA de novo sequencing, by reporting amino acid and peptide recalls and precisions compared with DIA-NN [27] search outcomes, considering only peptides unique to the test set. Our findings indicate that de novo peptide sequencing using DIA on older-generation instruments like the Q Exactive[28] or Fusion [29] is suboptimal, yielding relatively lower peptide recall. However, the Orbitrap Astral [30] significantly outpaces older-generation in acquisition speed, enabling the use of narrow-window data-independent acquisition (nDIA). The Astral narrow-window DIA method [30], which offers more consistent fragmentation patterns and fewer missing fragmentations, even in case of high coelution level, yields better results. Our extensive experiments across various datasets, including older-generation and Astral data, highlight the robustness and effectiveness of our proposed method. We demonstrated that our model consistently outperformed the baselines. These findings underscore the potential of our method for robust and accurate de novo peptide sequencing in the challenging DIA setting.
@@
To better understand these challenges, we developed a theoretical framework that explains how the balance between signal enhancement and noise accumulation in different acquisition methods affects de novo peptide sequencing performance. Specifically, we analyze the signal and noise characteristics of DDA, older-generation DIA, and Astral DIA, and their impact on the efficacy of peptide matching algorithms like XCorr.[31] [32] Our model reveals that older-generation DIA introduces substantial noise that outweighs the marginal gain in signal peaks—diminishing de novo sequencing performance—while Astral DIA provides a disproportionate increase in signal peaks despite higher noise levels. This increase effectively enhances peptide identification by improving the confidence of peptide-spectrum matches. This theoretical insight underscores the critical importance of optimizing signal-to-noise profiles in mass spectrometry data to improve de novo sequencing outcomes.
@@ 2 Results
Furthermore, in the comparison of DDA and DIA on the Hela [33] (older-generation) and PXD046453 [30] Hek293T (Astral) dataset, DIA demonstrated advantage in peptide detection when isolation window is small, with our de novo sequencing algorithm identifying additional peptides that DDA missed. These findings provide valuable insights into the proper scenarios for applying DIA de novo sequencing and highlight DIA’s capability to extend identification beyond DDA.
@@
We compared our model’s performance with the baselines across several datasets [34] [35] [33] in Figure 2, where peptide recall is defined as the proportion of peptides whose entire sequences exactly match the database search results, without any mismatches or errors. On most datasets, our model demonstrated significantly better performance than the baselines, however the results are relatively lower across the board compared to DDA levels [2] [14], indicating that there is a significant gap between database and de novo identification performance in older-generation mass spectrometers. Our amino acid recall is on average 60% higher than DeepNovo-DIA, while peptide recall being 53% higher. Comparing to Transformer DIA, our amino acid recall and peptide recall is 32% and 24% higher respectively.
@@
For older-generation instruments, we adopted the Hela dataset [33], analyzing peptides identified from the same biological sample in the DDA and DIA modes with varying DIA isolation windows (5 Th, 10 Th, and 20 Th). In the DDA experiments, we performed a database search using the PEAKS[36] DB search engine and employed PointNovo [37] for de novo peptide sequencing. For DIA, we utilized the DIA-NN[27] search engine for database search and applied our de novo sequencing methods.
@@
In contrast, the performance of the Orbitrap Astral mass spectrometer, using the PXD046453 [30] HEK293T dataset, presents a more consistent scenario (shown in Figure 5) due to the enabled narrow window mode. This dataset provides both DDA data and DIA data acquired at different cycle times (2.5 ms and 3.5 ms), both using a narrow 2 Th isolation window. The Astral data demonstrates that DIA consistently outperforms DDA with the narrow window mode enabled. The DDA de novo sequencing identifies approximately 17,000 peptides, whereas the DIA de novo sequencing detects nearly 32,000 peptides, showing a considerable increase over DDA. The additional 15,000 peptides identified by DIA de novo are likely missed by DDA due to biased sampling. This highlights the key advantage of DIA over DDA—its comprehensive and unbiased sampling of peptides, which is further enhanced by the speed and sensitivity of the Orbitrap Astral. In this case, the smaller isolation window and improved performance of Astral’s DIA mode enable the identification of a broader range of peptides, overcoming the limitations observed with DDA.
@@
Comparison with Cascadia [22]
@@
In database searches, the XCorr score [31] quantifies the similarity between the experimental spectrum and theoretical spectra generated from candidate peptides in the database. A higher XCorr score reflects a better match, leading to increased confidence in peptide identification. This is because the probability of a high-scoring match occurring by chance decreases exponentially with increasing score, enhancing the specificity of the identification. For de novo sequencing, the advantage of a higher XCorr score stems from the intrinsic similarities between database search algorithms and de novo methods. De novo algorithms are heuristic approaches designed to reconstruct peptide sequences directly from spectra without relying on a predefined database. Although de novo methods may employ advanced scoring functions tailored for sequence reconstruction, the XCorr score serves as a valuable lower-bound approximation of their performance. A higher XCorr score implies that the spectrum contains clear and informative fragment ion peaks that can be effectively utilized by de novo algorithms to deduce the peptide sequence.
@@
To validate our theory that signal-noise profile plays a crucial rule in de novo performance, we conducted simulations by performing a grid search through signal peak and noise peak numbers by generating synthetic spectra. Utilizing the approach proposed by Noble et al. (2012)[32], we computed the p-value for a single peptide-spectrum match based on dynamic programming and XCorr scoring. This method allows for the estimation of the statistical significance of the observed XCorr scores, providing insight into the confidence of peptide identifications under varying signal-to-noise conditions.
@@ 4 Methods
To better understand the statistical behavior of de novo peptide sequencing, we performed controlled simulations to generate synthetic spectra that mimic the signal and noise characteristics of different acquisition methods and test datasets. Signal peaks were sampled from theoretical spectra to represent peptide evidence, while noise peaks were added at levels representative of typical coelution and background noise for each acquisition method. For each simulated spectrum, we computed XCorr scores by cross-correlating the simulated spectra with the corresponding theoretical spectra, followed by significance estimation using a dynamic programming approach for p-value calculation [32]. Our analysis further examined the impact of multiple hypothesis testing in de novo sequencing, where the dramatically larger search space compared to database searches necessitates stricter statistical corrections, such as the Šidák adjustment [38]. This adjustment increases the stringency of significance thresholds, which in turn reduces peptide recall in de novo sequencing relative to database-driven methods. This theoretical derivation further illustrates that database searches are less impacted by the elevated p-values arising from the degraded signal-to-noise profiles observed when transitioning from DDA to older-generation DIA. This resilience helps database search methods maintain reasonable performance under noisier conditions, whereas de novo sequencing, with its vastly larger hypothesis space, experiences a more pronounced drop in performance.
@@
5 Conclusion
In this paper, we present a novel and highly effective method for de novo peptide sequencing in the DIA setting, offering significant improvements over traditional approaches. A key focus of our study is the comparison between DDA and DIA, where we demonstrate that while DIA outperforms DDA with narrow isolation windows, its advantage diminishes with wider windows in older-generation instruments. With Orbitrap Astral,, due to enabled narrow-window mode, DIA consistently surpasses DDA, highlighting the importance of next-generation mass spectrometers in improving peptide identification. To explain this phenomenon, we propose a simulation based theory, and highlight what aspects lead to improved identification performance.
Our method, specifically designed to tackle the highly mutliplexed DIA data, along with leveraging advanced instrumentation, provides robust performance in challenging proteomic environments. These findings emphasize the value of using DIA with modern technologies for comprehensive peptide detection, paving the way for more reliable and effective de novo sequencing methods in proteomics.
Acknowledgements
@@
Code Availability
The code for this project is available via https://github.com/hearthewind/dianovo.
@@ Appendix A The DIANovo Model
The feature construction for the encoder is designed to capture the chromatogram characteristics in DIA, in order to better distinguish signal peaks from coeluting fragment ions and noise. For each PSM (peptide-spectrum match), we collect five neighboring MS2 spectra along the retention time (RT) dimension, centered around the RT peak. The neighboring spectra provide a chromatogram for each observed peak in the spectrum, and the relationship among these chromatograms improves the robustness of peptide sequence predictions. [27]. During preprocessing, peaks across neighboring spectra with similar $m/z$ values are grouped together, enabling the construction of chromatograms for each m/z value. This is achieved by binning each spectrum into fixed-size 0.01 Th intervals, and once the chromatograms are generated through binning, the spectra are reconstructed as a list of (m/z, chromatogram) pairs for neural network processing. The binning approach offers greater computational efficiency than methods such as KD-trees, making it particularly well-suited for the characteristics of our DIA data.
@@
In addition to the primary spectral data, we also included eight supplementary features to enrich the feature set used for encoding, adapted from GraphNovo[14] and Novor[39]. These features are derived from each spectrum within the $\pm 50$ Th neighboring window and are designed to capture various statistical and intensity-based characteristics of the spectral peaks.
@@
where offset $\sigma$ depends on the ion type, computed as Supporting Table 1; $c$ is the charge of ion; $m_{proton}$ is the mass of proton and $m_{precursor}$ is the precursor mass [14].
@@
In addition, we adopt the FlashAttention 2 [40] implementation of the attention function, allowing us to process very large DIA spectra, with high computation and memory efficiency.
@@
GraphNovo employs the Graphformer[41] graph neural network to encode the spectrum graph, capturing comprehensive information about the spectrum. However, applying this approach to DIA data is impractical due to the large size of DIA spectra, which results from high levels of coelution. To address this challenge, we replace the memory-intensive node and path encoders used in Graphformer with Rotary Position Embedding (RoPE). This step is also neccessary for the adoption of Flash Attention 2, since it does not allow us to operate directly on the attention matrix, which is required by Graphformer.
RoPE is a position encoding technique that introduces rotational invariance by representing positional information in a circular format[42]. In our approach, RoPE encodes the mass differences between graph nodes, allowing the model to learn meaningful mass relationships automatically. This effectively transforms the spectrum graph into a fully-connected graph where edges represent mass differences, enabling the model to identify the most relevant mass differences for the task without the computational overhead of traditional node and path encoders. The adoption of RoPE is advantageous, as it captures the critical information conveyed by the mass difference between pairs of graph nodes, effectively representing the cumulative mass of the amino acids that lie between them.
@@
We adopt a two-stage decoding process, similar to GraphNovo[14] to alleviate the sequence memorization issue during training, where the decoder simply remembers seen training sequences, and fail to generalize on unseen sequences. This approach includes: